OpenStudy (josedavid):

please i need help please how can you solve the equation 4x = 2x + 2 graphically

OpenStudy (pinklion23):

hey :*

OpenStudy (pinklion23):

i know how to solve it but idk how to solve it graphically :(

OpenStudy (josedavid):

this is what i put Graph: y = 4x and y =2x + 2 Both will be straight lines. The point where they intersect will be the solution. also y = 4x y = 2x + 2 Plot that into desmos and you'll get (1,4) which is the solution 4 = 4(1) 4 = 2(1) + 2 4 = 4 they should intersect at (1,4) as we already determined from the table where when x = 1, both functions gave us 4.

OpenStudy (josedavid):

brb

OpenStudy (pinklion23):

http://www.algebra.com/algebra/homework/Graphs.faq.question.170807.html this might help

OpenStudy (josedavid):

i dont know is my answer is right